OpenStudy (darkigloo):

(calculus) The point on the curve x^2+2y = 0 that is nearest to the point (0,-1/2) occurs where y is?

myininaya (myininaya):

So you know we are talking about distance? And we want to minimize the distance?

OpenStudy (darkigloo):

ohh. ok. and how would i do that?

myininaya (myininaya):

do you know the distance formula

OpenStudy (darkigloo):

yes

myininaya (myininaya):

ok use that to find the distance between (x,y) and (0,-1/2) also if x^2+2y=0 then how can we rewrite y?

OpenStudy (darkigloo):

y= -x^2/2

myininaya (myininaya):

right now use the distance formula one (x,-x^2/2) and (0,-1/2)

OpenStudy (darkigloo):

\[\sqrt{-x^{2} + (-\frac{ 1 }{ 2 } + \frac{ x ^{2} }{ 2 }})^2\]

OpenStudy (darkigloo):

is that right?

myininaya (myininaya):

how did that - end up on the outside of that square?

OpenStudy (darkigloo):

i did (0 - x)^2

myininaya (myininaya):

which is x^2

myininaya (myininaya):

not -x^2

OpenStudy (darkigloo):

ohh right.

myininaya (myininaya):

you could say (-x)^2 but not -x^2

myininaya (myininaya):

but the square of (-x) is also the same as saying the square of x

myininaya (myininaya):

since the square of -1 is 1

myininaya (myininaya):

alright now you can find derivative of d^2 to find your critical numbers

myininaya (myininaya):

you could find it of d but d^2 is easier

OpenStudy (darkigloo):

so the derivative is x^3+x

myininaya (myininaya):

now set that equal to 0 to find critical numbers

OpenStudy (darkigloo):

x=0

myininaya (myininaya):

of we have the relationship between x and y is y=-x^2/2 so if x=0 then y=?

OpenStudy (darkigloo):

y=0

OpenStudy (darkigloo):

so 0 is the answer?

myininaya (myininaya):

yep that's what i got